Summarize key facts from this week's material. Create and solve a time value of money example and write a few paragraphs about how this could be useful in a corporate setting. You can follow examples from the homework and notes as your guide, but the problem you create should be original and you should provide the solution in Excel. The Reflection should be roughly one page in length
Answer #1

Summarize key facts from this week's material.

In this week we focused on time value of the money.

  • Once we have estimated the potential cash flows, we need to carry out a discounted cash flow analysis. The cash flows need to be discounted because there is a time value of money.
  • Cash flows which occur earlier are more valuable than cash flows which occur later.
  • The time value of money is important for three reasons: inflation, uncertainty, and opportunity costs. The discount rate by which cash flows are discounted should be good enough to compensate an investor for all the three parameters.
  • Time value discounts all the future cash flows of the firm with the required rate of return (RRR) to obtain their present values. Present value of a future cash flow is that value of cash required today that will grow to the future value if invested at RRR. Hence, present value of a cash flow (CFt) in period t (years) from now at a RRR of k is given by,

CFt PV(CFt) = (1 + k)

Example

Andrew Doughlas, an MBA from a well-known institute of management in western part of USA and currently the Head – Investments at XYZ Corporation is evaluating a capital budgeting proposal with cash flow profile shown below.The hurdle rate is 12%. Help Douglas calculate NPV of this project.

Time period (t)

Cash Flow ($ mn)

0

(500)

1

125

2

150

3

180

4

210

5

250

Solution

NPV = PV (Cash Inflows) – PV (Cash Outflows) = (1+ k)t - CF.

NPV = (500) + 125 / (1 + 12%)1 + 150 / (1 + 12%)2 + 180 / (1 + 12%)3 + 210 / (1 + 12%)4 + 250 / (1 + 12%)5 = (500) + 111.61 + 119.58 + 128.12 + 133.46 + 141.86 = $ 134.62 million

Otherwise the same can be calculated using a financial calculator or in excel using NPV or XNPV function. We can't provide you an excel fle, but i have provided the snapshot of my model below along with the excel formula used in the adjacent cell, so that you can create an excel file at your end easily.

1 Time period (t) Cash Flow ($ million) (500.00) 125.00 150.00 180.00 210.00 250.00 NPV 134.62 =NPV(12%,N3:N7)+N2

Thus in time values terms, the project is worth $ 134.62 million today. This is also called NPV. Since the project is NPV positive, this project must be accepted.